# DeepEN: Personalized Enteral Nutrition for Critically Ill Patients using Deep Reinforcement Learning

### v1

@@ Abstract
We introduce DeepEN, a deep reinforcement learning (RL) framework for personalized enteral nutrition (EN) in critically ill patients. Trained offline on over 11,000 ICU patients from the MIMIC-IV database, DeepEN generates 4-hourly recommendations for caloric, protein, and fluid intake tailored to each patient’s evolving physiology. The model integrates a curated, clinically informed state space with a custom reward function that balances short-term physiological and nutrition-related goals with long-term survival outcomes. Using a dueling double deep Q-network with conservative Q-learning regularization, DeepEN learns clinically realistic policies that align with high-value clinician actions while discouraging unsafe deviations. Across various qualitative and quantitative metrics, DeepEN outperforms clinician-derived and guideline-based policies, achieving a 3.7 ± 0.17 percentage-point reduction in estimated mortality (18.8% vs 22.5%) and improvements in key nutritional biomarkers. These findings highlight the potential of safe, data-driven personalization of EN therapy to improve outcomes beyond traditional guideline- or heuristic-based approaches.
@@ 1 Introduction
Enteral nutrition (EN), also known as tube feeding or enteral feeding, refers to the supply of nutrients such as carbohydrates, proteins, and fluids directly through the gastrointestinal tract. EN is typically initiated within 48 hours of Intensive Care Unit (ICU) admission for patients who are unable to meet their nutritional requirements through oral intake [1]. This can be due to increased metabolic demand or impairments in swallowing, nutrient absorption, or appetite regulation. Such conditions include mechanical ventilation, severe head injury, altered mental status, and gastrointestinal dysfunction [2]. The provision of adequate calories and protein is essential for such patients, but it remains clinically challenging.
@@
Evidence based guidelines such as from the American Society for Parenteral and Enteral Nutrition (ASPEN) provide weight and condition specific daily targets for calories and protein intake, considering factors such as obesity, renal replacement therapy, and hypermetabolic states [3]. However, these guidelines would continue to evolve as new clinical evidence emerge. Recent randomized controlled trials comparing aggressive verses conservative nutritional strategies in critically ill patients have reported conflicting results, ranging from improved outcomes to even potential harm [3, 4, 5]. These mixed findings underscore persistent uncertainties around the optimal dose and timing of enteral nutrition in the ICU setting.
@@
Moreover, there is a well-documented gap between guidelines and clinical practice. Surveys of ICU physicians and nurses reveal that lack of familiarity with both international and local nutrition guidelines, coupled with perceived gaps in the supporting evidence base, significantly hinders the consistent implementation of recommended enteral nutrition practices [6, 7, 8, 9] . Furthermore, nutritional decisions in the ICU are typically guided by heuristic-based practices or one-size-fits-all protocols, which fail to account for the dynamic and patient-specific trajectories of critically ill patients. Moreover factors such as evolving organ function, variable tolerance to feeding, competing interventions (e.g., vasopressors, sedation, dialysis) and rapidly varying metabolic demands make nutritional needs highly individualized and temporally variable. As a result, intervention inconsistencies such as underfeeding, overfeeding and delayed initiation remain common, contributing to increased infection risk, prolonged mechanical ventilation, muscle wasting, and mortality [10].
@@
Consequently, there is a pressing need address these limitations using adaptive, data-driven decision support systems that can effectively recommend personalized nutrition therapy in real time. Reinforcement learning (RL), which is a class of machine learning algorithms that learn optimal decision-making policies through interaction with complex environments, offers a promising framework for this task. RL models can learn from retrospective clinical data to recommend patient-specific, time-dependent feeding strategies that optimize long-term outcomes such as survival and recovery. Unlike static and population-wide guidelines, RL approaches can continuously adapt to a patient’s evolving physiological state, balance competing clinical goals, and offer individualized recommendation at each time point. While RL has shown promise in healthcare settings such as in sepsis [11, 12] and mechanical ventilation management [13], to our knowledge, it has yet to be utilized in the domain of enteral nutrition in critical care.
@@ 2.2 Q-learning and Offline RL
Q-learning [14] is one of the foundational algorithms in reinforcement learning (RL) and is among the most commonly used approaches in healthcare RL.[15] The method aims to learn the value of taking an action $a$ in a given state $s$, known as the Q-value $Q(s,a)$. At each time step $t$, the agent updates this estimate based on the reward $r_{t}$ received and the expected future return, using the Bellman update rule:
@@
To handle large or continuous state spaces, Deep Q-Networks (DQN) [16] approximate the Q-function using a neural network. While this leverages the expressive power of neural networks, it also introduces sensitivity to the distribution of training data. In particular, the presence of the maximization operator $\max_{a^{\prime}}Q(s_{t+1},a^{\prime})$ in the update rule can cause upward bias in the target, especially when Q-values are based on noisy or overfitted estimates, leading to overestimation errors.
This issue becomes more profound in offline (or batch) reinforcement learning, where the agent is limited to learning from a fixed dataset $\mathcal{D}$ collected by a behavior policy [17]. In healthcare RL, offline RL is the standard approach since direct interaction with the environment (i.e., directly acting on real patients) is unsafe and ethically impermissible. Without access to environmental interaction (as in online RL), the algorithm must generalize from the static dataset. Consequently, state-action pairs that are rare or missing in $\mathcal{D}$ may result in out-of-distribution (OOD) estimates, exacerbating overestimation error. In high-stakes domains like healthcare, such inaccurate Q-values can lead to unsafe or clinically inappropriate treatment recommendations if not properly mitigated.
@@ 2.3 Addressing overestimation bias with D3QN and CQL in offline RL
To mitigate the overestimation bias inherent in standard DQN, the Dueling Double DQN (D3QN) [18, 19, 20] algorithm was proposed as an enhancement. D3QN incorporates two key enhancements. First, it employs the Double Q-learning approach to decouple action selection and evaluation. The action is selected using the online network as $a^{\prime}=\arg\max_{a}Q(s_{t+1},a;\theta)$, and evaluates using a separate target network as $Q(s_{t+1},a^{\prime};\theta^{-})$. Second, D3QN incorporates a dueling network architecture that separately estimates the state-value function and the advantage function. This structure allows the agent to more effectively learn which states are valuable independently of the chosen actions, improving learning efficiency and stability. While these augmentations reduce the overestimation bias caused by the maximization operator over noisy Q-estimates, D3QN still does not fully resolve challenges in offline reinforcement learning, where distributional shift and insufficient coverage of state-action pairs remain significant limitations.
To further combat overestimation in offline reinforcement learning, Conservative Q-Learning (CQL) [21] was proposed. CQL addresses the risks posed by out-of-distribution (OOD) state-action pairs by explicitly penalizing overestimated Q-values for unseen or rarely seen actions. Specifically, CQL augments the Q-learning error with the term $\mathbb{E}_{s_{t}\sim\mathcal{D},a_{t}\sim\mathcal{A}}[Q(s_{t},a_{t})]$, which penalizes large Q-values for all possible actions at a given state, including those not sampled in the dataset. This discourages the agent from overvaluing out-of-distribution actions. In addition, CQL also includes the negative expectation $\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}}[Q(s_{t},a_{t})],$ which promotes higher Q-values for the actions that are actually observed in the dataset. Together, these terms ensure that the learned Q-function remains pessimistic about unsupported actions while maintaining high values for well-supported actions, implicitly steering the learned policy towards in-distribution (’clinician-like’) behavior. We combine both techniques of D3QN and CQL to avoid unsafe policy recommendations driven by overestimated Q-values in underrepresented areas of the state-action space.
@@ 2.4 RL for Enteral Nutrition Management
In the context of clinical decision support, RL has been widely explored for applications such as fluid and vasopressor dosing in sepsis patients, [11, 12], mechanical ventilation management, [13], and sedation titration[15]. However, to the best of our knowledge, RL has not yet been applied to the management of enteral nutrition (EN) in the ICU. Existing AI efforts in this domain have largely focused on predictive tools for malnutrition screening [22, 23] or detecting intolerance to EN [24, 25], rather than personalized nutritional support. Developing an RL-based nutrition management system presents both novel opportunities and unique challenges. Unlike pharmacologic or mechanical interventions (e.g., vasopressors or ventilator settings) that typically produce immediate and measurable physiological responses, nutritional interventions yield more gradual and diffuse effects, and often confounded by concurrent therapies and patient heterogeneity. This complicates the construction of a reward function that meaningfully encodes clinical outcomes such as mortality, recovery, or nutritional adequacy.
@@ 3 Methodology
Figure 1: Overview of solution pipeline.
@@ 3.1.1 States
Table 1: Table of State Space Variables
@@
Labs (24)* Albumin, pH, Calcium, Glucose, Hemoglobin, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, Lactate, Chloride, Platelets, Potassium, PTT, PT, AST, ALT, BUN, INR, Ionised Calcium, Total Bilirubin, Base Excess, , Phosphate
Feeding Related (5) Previous Calories, Previous Protein, Previous Water, Cumulative Sum of Calories, Cumulative Sum of Protein
Treatments and Interventions (8) Mechanical Ventilation, FiO2, CRRT, IV Fluids, Vasopressor Dose, Propofol Dose, Insulin Dose, 24-hour Cumulative Insulin Dose
@@
  • Abbreviations – CKD: Chronic Kidney Disease; DBP: Diastolic blood pressure; GCS: Glasgow Coma Scale; Hb: Hemoglobin; HR: Heart rate; ICU: Intensive care unit; MBP: Mean blood pressure; PTT: Partial Thromboplastin Time; Resp.: Respiratory rate; SBP: Systolic blood pressure; Temp.: Temperature; WBC: White blood cell
@@ 3.1.3 Rewards
We define a composite reward function that integrates both intermediate physiological indicators and long-term patient outcomes. The formulation is inspired by well-established reward structures used in healthcare RL research [11], and tailored to the clinical context of enteral nutrition.
@@
Consistent with established reward formulations in prior work in RL-based dynamic treatment regimes [11, 12], we set $r_{T}=15$, which preserves the dominance of the terminal outcome while remaining compatible with the scale of intermediate shaping signals.
@@
$R_{\text{phys}}=$ $\underbrace{-c_{0}\cdot\mathbf{1}[(\Delta\text{SOFA}=0)\land(\text{SOFA}_{t}>0)]}_{\text{Stagnation penalty}}+\underbrace{-c_{1}\cdot(\text{SOFA}_{t+1}-\text{SOFA}_{t})}_{\text{SOFA change penalty}}$
$+$ $\underbrace{-c_{2}\cdot\tanh(\text{Lactate}_{t+1}-\text{Lactate}_{t})}_{\text{Lactate worsening penalty}}$
@@
Intermediate Biomarker Deviation Reward ($R_{\text{bio}}$): We further incorporate two biomarkers — blood glucose and serum phosphate — due to their nutritional relevance and responsiveness to carbohydrate and protein intake [26, 27, 28, 29]. These biomarkers serve as strong signals indicative of feeding tolerance and general safety-related clinical outcomes [30, 31]. Interventions such as insulin, which can directly alter serum glucose levels, were included in the state space to account for their confounding effects on these biomarkers. The composite biomarker reward is defined as:
@@
For glucose, the target range is $[140,180]$ mg/dL [32], and for phosphate, the target range is $[2.5,4.5]$ mg/dL [33]. This formulation rewards both the absolute proximity of each biomarker to its target and any improvement over time.
@@ 3.2.1 Data Extraction and Preprocessing
We extracted a cohort of enteral feeding patients from the MIMIC-IV database [34] which contains data for over 60,000 ICU patients from the Beth Israel Deaconess Medical Center (Boston, MA, USA) from between 2008 to 2019. Inclusion criteria contained following conditions: 1) patient is >>18 years of age, 2) has at least 12 hours of enteral feeding data, and 3) has at least one recorded weight measurement. In total, data for 11,378 patients were included. For each patient, we collect data on demographics, vital signs, lab-values, fluids, specific diagnoses, enteral feeding data, and other relevant variables detailed in section 3.2. We only include data from periods where EN is administered, and did not consider periods where EN is absent or interrupted, nor did we consider data from parenteral feeding. In line with clinical evidence, we use enteral feeding data only from the first 10 days post-ICU admission to focus on the acute phase of critical illness (PCI) [35]. Data for calories, protein, and water intake were extracted from MIMIC-IV’s inputevents table. We aggregated the patient trajectories into four-hourly windows using mean or sum as appropriate, and fill any missing values using linear interpolation. More details on the patient cohort can be found in the Appendix (see Section 1: Cohort Information).

### v2

# DeepEN: A Deep Reinforcement Learning Framework for Personalized Enteral Nutrition in Critical Care

@@ Abstract
Objective: Current ICU enteral feeding remains sub-optimal due to limited personalization and ongoing uncertainty about appropriate calorie, protein, and fluid targets—particularly in the context of rapidly changing metabolic demands and heterogeneous responses to therapeutic interventions. This study introduces DeepEN, a novel reinforcement learning (RL)–based framework designed to dynamically personalize enteral nutrition (EN) dosing for critically ill patients using electronic health record data.
Methods: DeepEN was trained on data from over 11,000 ICU patients in the MIMIC-IV database to generate 4-hourly, patient-specific targets for caloric, protein, and fluid intake. The model’s state space integrates demographics, comorbidities, vital signs, laboratory measurements, and recent interventions considered relevant to nutritional management. The reward function was designed with domain expertise to balance short-term physiological and nutrition-related goals with long-term survival outcomes, reflecting real-world clinical priorities. The framework employs a dueling double deep Q-network with Conservative Q-Learning regularization to ensure safe and reliable policy learning from retrospective data. Model performance was benchmarked against both clinician-derived and guideline-based policies.
Results: DeepEN outperformed both clinician and guideline-based policies, achieving a 3.7 ± 0.17 percentage-point absolute reduction in estimated mortality compared with the clinician policy (18.8% vs 22.5%) and higher expected returns relative to the gold-standard guideline policy (11.89 vs 8.11). Control of key nutritional biomarkers was also improved under the learned policy. U-shaped associations were observed between deviations from clinician dosing and mortality for caloric, protein, and fluid recommendations, suggesting that the learned policy aligns with clinician actions associated with higher survival while diverging from suboptimal behaviors.
Conclusion: DeepEN demonstrates the feasibility and potential of conservative offline RL for safe, individualized enteral nutrition therapy in critical care. These findings suggest that data-driven personalization may enhance patient outcomes beyond traditional guideline- or heuristic-based approaches.
keywords:
Reinforcement learning , critical care , enteral nutrition , personalized medicine
Jointly supervised this work.
*Corresponding author: ephfm@nus.edu.sg
@@ 1 Introduction
Enteral nutrition (EN), also known as tube feeding or enteral feeding, refers to the supply of nutrients such as carbohydrates, proteins, and fluids directly through the gastrointestinal tract. EN is typically initiated within 48 hours of Intensive Care Unit (ICU) admission for patients who are unable to meet their nutritional requirements through oral intake [adeyinka2018enteric]. This can be due to increased metabolic demand or impairments in swallowing, nutrient absorption, or appetite regulation. Such conditions include mechanical ventilation, severe head injury, altered mental status, and gastrointestinal dysfunction [preiser2021guide]. The provision of adequate calories and protein is essential for such patients, but it remains clinically challenging.
@@
Evidence based guidelines such as from the American Society for Parenteral and Enteral Nutrition (ASPEN) provide weight and condition specific daily targets for calories and protein intake, considering factors such as obesity, renal replacement therapy, and hypermetabolic states [compher2022guidelines]. However, these guidelines would continue to evolve as new clinical evidence emerge. Recent randomized controlled trials comparing aggressive verses conservative nutritional strategies in critically ill patients have reported conflicting results, ranging from improved outcomes to even potential harm [compher2022guidelines, bels2024effect, heyland2023effect]. These mixed findings underscore persistent uncertainties around the optimal dose and timing of enteral nutrition in the ICU setting.
@@
Moreover, there is a well-documented gap between guidelines and clinical practice. Surveys of ICU physicians and nurses reveal that lack of familiarity with both international and local nutrition guidelines, coupled with perceived gaps in the supporting evidence base, significantly hinders the consistent implementation of recommended enteral nutrition practices [mirhosiny2021physicians, friesecke2014improvement, jarden2015practice, ellis2015improving] . Furthermore, nutritional decisions in the ICU are typically guided by heuristic-based practices or one-size-fits-all protocols, which fail to account for the dynamic and patient-specific trajectories of critically ill patients. Moreover factors such as evolving organ function, variable tolerance to feeding, competing interventions (e.g., vasopressors, sedation, dialysis) and rapidly varying metabolic demands make nutritional needs highly individualized and temporally variable. As a result, intervention inconsistencies such as underfeeding, overfeeding and delayed initiation remain common, contributing to increased infection risk, prolonged mechanical ventilation, muscle wasting, and mortality [ramaswamy2024nine].
@@
Consequently, there is a pressing need address these limitations using adaptive, data-driven decision support systems that can effectively recommend personalized nutrition therapy in real time. Reinforcement learning (RL), which is a class of machine learning algorithms that learn optimal decision-making policies through interaction with complex environments, offers a promising framework for this task. RL models can learn from retrospective clinical data to recommend patient-specific, time-dependent feeding strategies that optimize long-term outcomes such as survival and recovery. Unlike static and population-wide guidelines, RL approaches can continuously adapt to a patient’s evolving physiological state, balance competing clinical goals, and offer individualized recommendation at each time point. While RL has shown promise in healthcare settings such as in sepsis [raghu2017deep, tan2024advancing] and mechanical ventilation management [peine2021development], to our knowledge, it has yet to be utilized in the domain of enteral nutrition in critical care.
@@ 2.2 Q-learning and Offline RL
Q-learning [watkins1992q] is one of the foundational algorithms in reinforcement learning (RL) and is among the most commonly used approaches in healthcare RL. [yu2021reinforcement] The method aims to learn the value of taking an action $a$ in a given state $s$, known as the Q-value $Q(s,a)$. At each time step $t$, the agent updates this estimate based on the reward $r_{t}$ received and the expected future return, using the Bellman update rule:
@@
To handle large or continuous state spaces, Deep Q-Networks (DQN) [mnih2015human] approximate the Q-function using a neural network. While this leverages the expressive power of neural networks, it also introduces sensitivity to the distribution of training data. In particular, the presence of the maximization operator $\max_{a^{\prime}}Q(s_{t+1},a^{\prime})$ in the update rule can cause upward bias in the target, especially when Q-values are based on noisy or overfitted estimates, leading to overestimation errors.
This issue becomes more profound in offline (or batch) reinforcement learning, where the agent is limited to learning from a fixed dataset $\mathcal{D}$ collected by a behavior policy [levine2020offline]. In healthcare RL, offline RL is the standard approach since direct interaction with the environment (i.e., directly acting on real patients) is unsafe and ethically impermissible. Without access to environmental interaction (as in online RL), the algorithm must generalize from the static dataset. Consequently, state-action pairs that are rare or missing in $\mathcal{D}$ may result in out-of-distribution (OOD) estimates, exacerbating overestimation error. In high-stakes domains like healthcare, such inaccurate Q-values can lead to unsafe or clinically inappropriate treatment recommendations if not properly mitigated.
@@ 2.3 Addressing overestimation bias with D3QN and CQL in offline RL
To mitigate the overestimation bias inherent in standard DQN, the Dueling Double DQN (D3QN) [van2016deep, wang2016dueling, hessel2018rainbow] algorithm was proposed as an enhancement. D3QN incorporates two key enhancements. First, it employs the Double Q-learning approach to decouple action selection and evaluation. The action is selected using the online network as $a^{\prime}=\arg\max_{a}Q(s_{t+1},a;\theta)$, and evaluates using a separate target network as $Q(s_{t+1},a^{\prime};\theta^{-})$. Second, D3QN incorporates a dueling network architecture that separately estimates the state-value function and the advantage function. This structure allows the agent to more effectively learn which states are valuable independently of the chosen actions, improving learning efficiency and stability. While these augmentations reduce the overestimation bias caused by the maximization operator over noisy Q-estimates, D3QN still does not fully resolve challenges in offline reinforcement learning, where distributional shift and insufficient coverage of state-action pairs remain significant limitations.
To further combat overestimation in offline reinforcement learning, Conservative Q-Learning (CQL) [kumar2020conservative] was proposed. CQL addresses the risks posed by out-of-distribution (OOD) state-action pairs by explicitly penalizing overestimated Q-values for unseen or rarely seen actions. Specifically, CQL augments the Q-learning error with the term $\mathbb{E}_{s_{t}\sim\mathcal{D},a_{t}\sim\mathcal{A}}[Q(s_{t},a_{t})]$, which penalizes large Q-values for all possible actions at a given state, including those not sampled in the dataset. This discourages the agent from overvaluing out-of-distribution actions. In addition, CQL also includes the negative expectation $\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}}[Q(s_{t},a_{t})],$ which promotes higher Q-values for the actions that are actually observed in the dataset. Together, these terms ensure that the learned Q-function remains pessimistic about unsupported actions while maintaining high values for well-supported actions, implicitly steering the learned policy towards in-distribution (’clinician-like’) behavior. We combine both techniques of D3QN and CQL to avoid unsafe policy recommendations driven by overestimated Q-values in underrepresented areas of the state-action space.
@@ 2.4 RL for Enteral Nutrition Management
In the context of clinical decision support, RL has been widely explored for applications such as fluid and vasopressor dosing in sepsis patients, [raghu2017deep, tan2024advancing], mechanical ventilation management, [peine2021development], and sedation titration [yu2021reinforcement]. However, to the best of our knowledge, RL has not yet been applied to the management of enteral nutrition (EN) in the ICU. Existing AI efforts in this domain have largely focused on predictive tools for malnutrition screening [besculides2023implementing, raphaeli2023using] or detecting intolerance to EN [hu2022development, chen2023development], rather than personalized nutritional support. Developing an RL-based nutrition management system presents both novel opportunities and unique challenges. Unlike pharmacologic or mechanical interventions (e.g., vasopressors or ventilator settings) that typically produce immediate and measurable physiological responses, nutritional interventions yield more gradual and diffuse effects, and often confounded by concurrent therapies and patient heterogeneity. This complicates the construction of a reward function that meaningfully encodes clinical outcomes such as mortality, recovery, or nutritional adequacy.
@@ 3 Methodology
Figure 1: Overview of the DeepEN solution pipeline.
@@ 3.1.1 States
Table 1: State Space Variables
@@
Labs (24)* Albumin, pH, Calcium, Glucose, Hemoglobin, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, Lactate, Chloride, Platelets, Potassium, PTT, PT, AST, ALT, BUN, INR, Ionised calcium, Total bilirubin, Base excess, Phosphate
Feeding Related (5) Previous calories, Previous protein, Previous water, Cumulative calories, Cumulative protein
Treatments and Interventions (8) Mechanical ventilation, FiO2, CRRT, IV fluids, Vasopressor dose, Propofol dose, Insulin dose, 24-h cumulative insulin dose
@@
Abbreviations: CKD = chronic kidney disease; DBP = diastolic blood pressure; GCS = Glasgow Coma Scale; HR = heart rate; ICU = intensive care unit; MBP = mean blood pressure; PTT = partial thromboplastin time; Resp. = respiratory rate; SBP = systolic blood pressure; WBC = white blood cell.
@@ 3.1.3 Rewards
We define a composite reward function that integrates both intermediate physiological indicators and long-term patient outcomes. The formulation is inspired by well-established reward structures used in healthcare RL research [raghu2017deep], and tailored to the clinical context of enteral nutrition.
@@
Consistent with established reward formulations in prior work in RL-based dynamic treatment regimes [raghu2017deep, tan2024advancing], we set $r_{T}=15$, which preserves the dominance of the terminal outcome while remaining compatible with the scale of intermediate shaping signals.
@@
$R_{\mathrm{phys}}={}$ $-c_{0}\mathbf{1}\{\,\Delta\mathrm{SOFA}=0\ \land\ \mathrm{SOFA}_{t}>0\,\}$ (stagnation) (1)
$-c_{1}(\mathrm{SOFA}_{t+1}-\mathrm{SOFA}_{t})$ (SOFA change)
$-c_{2}\,\tanh(\mathrm{Lactate}_{t+1}-\mathrm{Lactate}_{t})$ (lactate penalty)
@@
Intermediate Biomarker Deviation Reward ($R_{\text{bio}}$): We further incorporate two biomarkers — blood glucose and serum phosphate — due to their nutritional relevance and responsiveness to carbohydrate and protein intake [gosmanov2013management, uribarri2007phosphorus, kalantar2010understanding, ikizler2013prevention]. These biomarkers serve as strong signals indicative of feeding tolerance and general safety-related clinical outcomes [mehanna2008refeeding, nguyen2007relationship]. Interventions such as insulin, which can directly alter serum glucose levels, were included in the state space to account for their confounding effects on these biomarkers. The composite biomarker reward is defined as:
@@
For glucose, the target range is $[140,180]$ mg/dL [american202416], and for phosphate, the target range is $[2.5,4.5]$ mg/dL [geerse2010treatment]. This formulation rewards both the absolute proximity of each biomarker to its target and any improvement over time.
@@ 3.2.1 Data Extraction and Preprocessing
We extracted a cohort of enteral feeding patients from the MIMIC-IV database [johnson2023mimic] which contains data for over 60,000 ICU patients from the Beth Israel Deaconess Medical Center (Boston, MA, USA) from between 2008 to 2019. Inclusion criteria contained following conditions: 1) patient is >>18 years of age, 2) has at least 12 hours of enteral feeding data, and 3) has at least one recorded weight measurement. In total, data for 11,378 patients were included. For each patient, we collect data on demographics, vital signs, lab-values, fluids, specific diagnoses, enteral feeding data, and other relevant variables detailed in section 3.2. We only include data from periods where EN is administered, and did not consider periods where EN is absent or interrupted, nor did we consider data from parenteral feeding. In line with clinical evidence, we use enteral feeding data only from the first 10 days post-ICU admission to focus on the acute phase of critical illness (PCI) [iwashyna2015towards]. Data for calories, protein, and water intake were extracted from MIMIC-IV’s inputevents table. We aggregated the patient trajectories into four-hourly windows using mean or sum as appropriate, and fill any missing values using linear interpolation. More details on the patient cohort can be found in the Appendix (see Section 1: Cohort Information).